A toroid core has N = 1200 turns, length L = 80cm, cross-sectional area A = 60cm2



,



current I = 1.5A. Compute B and H. Assume an empty core.

1
Expert's answer
2022-03-21T12:43:37-0400

Answer

Magnetic field due to toroid

"B=\frac{\mu_0Ni}{2\pi r}\\=\frac{2*10^{-7}*1.5}{0.044}\\=68.64*10^{-7}T"

Now H

"H=\frac{B}{\mu_0}\\=\frac{68.64*10^{-7}}{4*3.14*10^{-7}}\\=5.47T"